This is from a pre-algebra worksheet. We do not understand how to solve for P when there is a fraction contained in a fraction. Thanks for any help you can give!
2 1/2 P
----- = -----
5 4
Read as: 2 1/2 over 5 equals P over 4. Solve for P

Answer by pleasuretohelp(9)

2 1/2 P
______ = _______
5 4
You can actually cross-multiplied to get rid of the fraction.
What I mean by cross-multiplying is:
a = c
__ ___
b d
(read: a over b equals c over d)
becomes: a * d = b * c or d * a = c * b. Either way, they are the same.
So, to answer your question:
(2 1/2) * 4 = 5 * P
(5/2) * 4 = 5 * P
20/2 = 5 * P
10 = 5 * P
Divide both sides by 5 in order to get P:
10/5 = (5 * P)/5
2 = P
P = 2